OpenStudy (anonymous):

How can I solve this indefinite integral without having arcsin in the result? \int \sqrt{1-x^2} dx Thanks!

OpenStudy (anonymous):

The integral is this: \[\int\limits \sqrt{1-x^2} dx\]

OpenStudy (anonymous):

i dont see how you got arcsin? /int for arsin is 1/(a^2 -x^2)^.5

OpenStudy (anonymous):

\[1/\sqrt{(a^2 - x^2)}\]

OpenStudy (anonymous):

I solved it with this substitution x=sin(t) result: 1/2arcsin(x)+1/4sin(2arcsin(x))

OpenStudy (anonymous):

I LOL YOU